What 3 numbers all less than 10 have their sum equal to there product
Answer by fcabanski(1391)

If they have to be distinct (different) positive whole numbers, use trial and error.


0*1*2 = 0+1+2? No.


1*2*3 = 1+2+3 = 6. Yes.